you invest $350 that pays a yearly interest rate of 6%
write a formula for an exponential function giving the balance of the account (B) , in dollars , as a function of the time (t) , in years, since initial investment

Answer by Theo(2967)

350 * (1.06)^t = future value of this account.

at the end of 1 year, the account will have 350 * (1.06)^1 = 371

at the end of 2 years, the account will have 350 * (1.06)2 = 393.26

that comes out as 350 * 1.06 * 1.06

since 350 * 1.06 = 371 * 1.06 = 393.26, the interest earned each year is combined with the principal and then interest is earned on the sum.

at the end of 10 years, the account will have 350 * (1.06)^10 = 626.80